OpenStudy (anonymous):

well, how to show 1/[2 X root 2 ] as an exponent of "2" ? GCSE Systems suck!

OpenStudy (saifoo.khan):

GCE RULES. :D

OpenStudy (anonymous):

this is not an answer!!!!!!!!!

OpenStudy (saifoo.khan):

\[\frac{1}{2\sqrt{x}}^2\] Like that?

OpenStudy (anonymous):

\[1\div(2\sqrt{2})^{2}\]

OpenStudy (saifoo.khan):